OpenStudy (anonymous):

a sample of 148 firecrackers contained 6 duds. how many duds would you expect in a sample of 1924 firecrackers

OpenStudy (anonymous):

same like previous, unit rule is used 148-----6 1-----6/148 1924------6*1924/148=78

OpenStudy (anonymous):

6 duds / 148 firecrackers X duds / 1924 firecrackers 6/148 = x/1924 148x=6*1924 148x=11544 x=78 Darnit he beat me to it.